Yoko invested her savings into investment funds. The $24,000 that she invested in fund A returned a 2% profit the amount that she invested in fund B returned a 6% profit. How much did she invest in fund B, if both funds together return the 3% profit
Answer by stanbon(75887)

Yoko invested her savings into investment funds. The $24,000 that she invested in fund A returned a 2% profit the amount that she invested in fund B returned a 6% profit. How much did she invest in fund B, if both funds together return the 3% profit
--------
interest + interest = interest
0.02*24000 + 0.06B = 0.03(24000+B)
-----------------------------------------
2*24000 + 6B = 3*24000 + 3B
-----
3B = 24000
B = $8000 (Amt of B invested)
------------
Cheers,
Stan H.
------------